# Motion Control in Multi-Rotor Aerial Robots Using Deep Reinforcement Learning

### v1

@@ Abstract
This paper investigates the application of Deep Reinforcement Learning (DRL) to address motion control challenges in drones for additive manufacturing (AM). Drone-based additive manufacturing promises flexible and autonomous material deposition in large-scale or hazardous environments. However, achieving robust real-time control of a multi-rotor aerial robot under varying payloads and potential disturbances remains challenging. Traditional controllers like PID often require frequent parameter re-tuning, limiting their applicability in dynamic scenarios. We propose a DRL framework that learns adaptable control policies for multi-rotor drones performing waypoint navigation in AM tasks. We compare Deep Deterministic Policy Gradient (DDPG) and Twin Delayed Deep Deterministic Policy Gradient (TD3) within a curriculum learning scheme designed to handle increasing complexity. Our experiments show TD3 consistently balances training stability, accuracy and success, particularly when mass variability is introduced. These findings provide a scalable path toward robust, autonomous drone control in additive manufacturing.
@@ I Introduction
However, the deployment of UAVs for AM tasks presents control challenges [4]. As drones deposit materials, their mass distribution changes continuously, altering both the center of gravity and inertia. In addition, external disturbances further complicate control, making it difficult to maintain stability and accuracy in dynamic three-dimensional environments [5]. These factors necessitate the development of robust and adaptive control systems.
@@
Reinforcement Learning (RL) offers an alternative approach to UAV control by enabling systems to learn control policies through interaction with their environment, without depending on explicit system models. This model-free method allows RL algorithms to adapt dynamically to both internal variations and external disturbances [10, 11]. However, standard RL methods encounter challenges in high-dimensional environments like AM, where state and action spaces are extensive and the environment is highly unpredictable [12].
@@
A key consideration in applying DRL to UAV-based AM is the choice between deterministic and stochastic policy methods. Deterministic approaches, such as Deep Deterministic Policy Gradient (DDPG) [14], produce specific actions for a given state. This ensures high precision and consistency [15], which are essential for accurate material deposition in AM. Stochastic methods sample actions from probability distributions, introducing variability that can reduce precision. Deterministic methods are better suited for UAV-based AM as they provide predictable and repeatable control. They also reduce computational overhead and work well in stable, low-noise environments [16]. However, DDPG faces challenges such as instability and overestimation bias [11]. Twin Delayed Deep Deterministic Policy Gradient (TD3) [17] addresses these issues. It uses twin critics to reduce overestimation and delayed policy updates to improve stability. These improvements make TD3 an effective choice for UAV-based AM tasks that require precise and reliable control under varying conditions [18].
@@ II Problem Formulation And Preliminaries
where $m$ is the UAV mass. This additional acceleration is incorporated into the UAV’s motion model in the simulation to effectively account for the impact of material deposition. Additionally, random noise is introduced into sensor measurements to ensure robustness against real-world uncertainties. The environment is assumed to be obstacle-free, allowing the study to focus solely on the effects of waypoint tracking, external disturbances, and system variations. The UAV is controlled through roll, pitch, and thrust, with yaw control omitted to reduce complexity while preserving stability. This design choice is justified as AM applications primarily rely on precise lateral and vertical positioning rather than continuous yaw adjustments.
@@ III-A System Architecture
The RL agent controls the multirotor using three actions: roll ($\phi_{a}$), pitch ($\theta_{a}$), and thrust ($T$). Yaw control is omitted in this configuration. Excluding yaw control reduces the complexity of the problem and allows the RL agent to stabilize and control the roll and pitch axes effectively. These actions are normalized to ensure compatibility with neural network policies and facilitate robust learning.
@@
The transition function $P(s_{t+1}\mid s_{t},a_{t})$ delineates how the drone transitions from its current state $s_{t}$ to the next state $s_{t+1}$ upon executing action $a_{t}$ within the RL framework. In this study, the transition dynamics are modeled deterministically using MATLAB Simulink’s UAV Toolbox, which simulates the multi-rotor’s physical behavior based on applied control inputs. The high-level actions are first scaled to match the UAV’s operational ranges and then translated into low-level Pulse-Width Modulation (PWM) [26] signals via a state estimator. The PWM commands directly influence the drone’s actuators as governed by the UAV’s dynamic model within Simulink. Consequently, the next state $s_{t+1}$ is deterministically determined by the interplay between the current state, the scaled actions, and the inherent physical laws modeled in the simulation environment.
The episode terminates when the drone reaches the target position with near-zero velocity. Conversely, termination also occurs under failure scenarios, if the drone crashes by touching the ground or if it deviates too far from the target position.
III-B Algorithm Evaluation
@@
III-C Curriculum Learning for Complexity Management
@@ IV Experiments
This paper employs MATLAB’s Simulink platform, along with the RL Toolbox and UAV Toolbox, to develop a simulation environment for a drone engaged in AM. The Simulink setup includes essential blocks such as the RL agent, observation, reward, and termination blocks, designed to enable real-time adaptive learning for stable control. The architecture, shown in Fig 2, outlines the interconnected components that drive the simulation and learning processes.
@@
In addition, curriculum learning is employed to further expedite the agent’s training process and improve convergence by gradually increasing the complexity of tasks as the agent’s proficiency improves. Table I summarizes the hyperparameter settings for the DDPG and TD3 agents.
@@
In addition, each critic network is initialized using a weight scaling strategy, where weights are sampled from a uniform distribution within the range
@@ V-A Performance on Target-Reaching Task
TD3 703 133.213 0.01804 0.000226 96
DDPG 604 140.603 0.05643 0.006412 82
@@ V-C Adaptation to Dynamic Mass Changes
Retraining the agent with the expanded observation space led to significant improvements. The average cumulative reward increased to $760$ and the success ratio rose to $94\%$ with a reduced average positional error of $0.09857m$ and a better precision of 0.0626.
@@ V-D Summary of Agent Performance
DDPG Single Waypoint Navigation No Noise Added; No Variable Mass 604 140.6029 0.05643 0.006412 82
TD3 Single Waypoint Navigation No Noise Added; No Variable Mass 703 133.2130 0.01804 0.000226 96
TD3 (No Curriculum Learning) Multiple Waypoint Navigation Noise Added; No Variable Mass 593 145.9763 0.01994 0.002780 66
TD3 (With Curriculum Learning) Multiple Waypoint Navigation Noise Added; No Variable Mass 729 148.5167 0.01581 0.006831 87
TD3 (With Acceleration in Observation Space) Multiple Waypoint Navigation Noise Added; Variable Mass 760 152.5167 0.09857 0.062600 94
@@ VI Conclusion
This study evaluated the performance of two RL algorithms, DDPG and TD3, for UAV control in additive manufacturing tasks. Results showed that TD3 outperformed DDPG in key metrics, including cumulative reward, average positional error, precision, and success ratio, confirming its suitability for precise and consistent target-reaching tasks.

### v2

@@ Abstract
This paper investigates the application of Deep Reinforcement Learning (DRL) to address motion control challenges in drones for additive manufacturing (AM). Drone-based additive manufacturing offers a flexible and autonomous solution for material deposition in large-scale or hazardous environments. However, achieving robust real-time control of a multi-rotor aerial robot under varying payloads and potential disturbances remains challenging. Traditional controllers like PID often require frequent parameter re-tuning, limiting their applicability in dynamic scenarios. We propose a DRL framework that learns adaptable control policies for multi-rotor drones performing waypoint navigation in AM tasks. We compare Deep Deterministic Policy Gradient (DDPG) and Twin Delayed Deep Deterministic Policy Gradient (TD3) within a curriculum learning scheme designed to handle increasing complexity. Our experiments show TD3 consistently balances training stability, accuracy and success, particularly when mass variability is introduced. These findings provide a scalable path toward robust, autonomous drone control in additive manufacturing.
@@ I Introduction
However, the deployment of UAVs for AM tasks presents control challenges [4]. As drones deposit materials, their mass distribution changes continuously, altering both the center of gravity and inertia. In addition, external disturbances further complicate control, making it difficult to maintain stability and accuracy in dynamic three-dimensional environments [5]. These challenges highlight the need for robust and adaptive control systems [6].
@@
Reinforcement Learning (RL) offers an alternative approach to UAV control by enabling systems to learn control policies through interaction with their environment, without depending on explicit system models. This model-free method allows RL algorithms to adapt dynamically to both internal variations and external disturbances [12]. However, standard RL methods encounter challenges in high-dimensional environments like AM, where state and action spaces are extensive and the environment is highly unpredictable [14, 13].
@@
A key consideration in applying DRL to UAV-based AM is the choice between deterministic and stochastic policy methods. Deterministic approaches, such as Deep Deterministic Policy Gradient (DDPG) [16], produce specific actions for a given state. This ensures higher precision and consistency [17], which are essential for accurate material deposition in AM. Stochastic methods sample actions from probability distributions, introducing variability that can reduce precision. Deterministic methods are better suited for UAV-based AM as they provide predictable and repeatable control. They also reduce computational overhead and work well in stable, low-noise environments [18]. However, DDPG faces challenges such as instability and overestimation bias [12]. Twin Delayed Deep Deterministic Policy Gradient (TD3) [19] addresses these issues. It uses twin critics to reduce overestimation and delayed policy updates to improve stability. These improvements make TD3 an effective choice for UAV-based AM tasks that require precise and reliable control under varying conditions [20].
@@ II Problem Formulation And Preliminaries
where $m$ is the UAV mass. This additional acceleration is incorporated into the UAV’s motion model in the simulation to effectively account for the impact of material deposition. Additionally, random noise is introduced into sensor measurements to ensure robustness against real-world uncertainties. A uniform noise model was used, with only measurement noise (no process noise), and this noise was injected into the state space of the drone once we obtain it from the multirotor model in the simulation. The environment is assumed to be obstacle-free, allowing the study to focus solely on the effects of waypoint tracking, external disturbances, and system variations. The UAV is controlled through roll, pitch, and thrust, with yaw control omitted to reduce complexity while preserving stability. The rotational dynamics, including attitude control, are managed separately using the MATLAB UAV Toolbox multirotor model (version 2024a), which ensures stability by converting control inputs to PWM commands. This design choice is justified as AM applications primarily rely on precise lateral and vertical positioning rather than continuous yaw adjustments.
@@ III-A System Architecture
The RL agent controls the multirotor using three actions: roll ($\phi_{a}$), pitch ($\theta_{a}$), and thrust ($T$). Yaw control is omitted in this configuration. Excluding yaw control reduces the complexity of the problem and allows the RL agent to stabilize and control the roll and pitch axes effectively. These actions are normalized to ensure compatibility with neural network policies and facilitate robust learning. The rotational dynamics, including attitude control, are managed by a separate algorithm using the MATLAB UAV Toolbox multirotor model (version 2024a), which handles the attitude rate control and converts the control inputs to PWM commands for stability.
@@
The transition function $P(s_{t+1}\mid s_{t},a_{t})$ delineates how the drone transitions from its current state $s_{t}$ to the next state $s_{t+1}$ upon executing action $a_{t}$ within the RL framework. In this study, the transition dynamics are modeled deterministically using MATLAB Simulink’s UAV Toolbox (version 2024a), which simulates the multi-rotor’s physical behavior based on applied control inputs. The high-level actions are first scaled to match the UAV’s operational ranges and then translated into low-level Pulse-Width Modulation (PWM) [28] signals via a state estimator. The PWM commands directly influence the drone’s actuators as governed by the UAV’s dynamic model within Simulink. Consequently, the next state $s_{t+1}$ is deterministically determined by the interplay between the current state, the scaled actions, and the inherent physical laws modeled in the simulation environment.
The episode terminates when the drone reaches the target position with a velocity less than $0.1m/s$ (near-zero velocity). Conversely, termination occurs under failure scenarios if the drone crashes by touching the ground (defined as a height less than 0.1 meters) or if it deviates more than 10 meters from the target position.
@@
III-B Curriculum Learning for Complexity Management
@@ IV Experiments
This paper employs MATLAB’s Simulink platform (version 2024a), along with the RL Toolbox and UAV Toolbox, to develop a simulation environment for a drone engaged in AM. The Simulink setup shown in Fig 2 includes essential blocks such as the RL agent, observation, reward, and termination function, designed to enable real-time adaptive learning for stable control.
@@
Table I summarizes the hyperparameter settings for the DDPG and TD3 agents.
@@
In addition, each critic network is initialized using a weight scaling strategy [32], where weights are sampled from a uniform distribution within the range
@@ V Results
In the test phase, the drone always started from a fixed, stationary position, and the waypoints it needed to navigate to changed with each trial. In contrast, during training, both the drone’s starting position and the waypoints were initialized randomly to expose the agent to a more diverse set of conditions. This approach helped the agent to become more robust by training in a variety of environmental configurations, ensuring its generalization ability across different scenarios.
@@ V-A Performance on Target-Reaching Task
TD3 703 133.213 $1.804\times 10^{-2}$ $2.26\times 10^{-4}$ 96
DDPG 604 140.603 $5.643\times 10^{-2}$ $6.412\times 10^{-3}$ 82
@@ V-C Adaptation to Dynamic Mass Changes
Retraining the agent with the expanded observation space led to significant improvements. The average cumulative reward increased from $454$ to $760$, and the success ratio rose from $29\%$ to $94\%$. Additionally, the average positional error was reduced from $0.2688m$ to $0.09857m$, and the precision (standard deviation of positional error), improved from $0.1459$ to $0.0626$.
@@ V-D Summary of Agent Performance
DDPG Single Waypoint Navigation No Noise Added; No Variable Mass 604 140.6029 $5.643\times 10^{-2}$ $6.412\times 10^{-3}$ 82
TD3 Single Waypoint Navigation No Noise Added; No Variable Mass 703 133.2130 $1.804\times 10^{-2}$ $2.26\times 10^{-4}$ 96
TD3 (No Curriculum Learning) Multiple Waypoint Navigation Noise Added; No Variable Mass 593 145.9763 $1.994\times 10^{-2}$ $2.780\times 10^{-3}$ 66
TD3 (With Curriculum Learning) Multiple Waypoint Navigation Noise Added; No Variable Mass 729 148.5167 $1.581\times 10^{-2}$ $6.831\times 10^{-3}$ 87
TD3 (With Acceleration in Observation Space) Multiple Waypoint Navigation Noise Added; Variable Mass 760 152.5167 $9.857\times 10^{-2}$ $6.260\times 10^{-2}$ 94
@@ VI Conclusion
This study evaluated the performance of two RL algorithms, DDPG and TD3, for UAV control in additive manufacturing tasks. Results showed that TD3 outperformed DDPG in key metrics, including cumulative reward, average positional error, precision, and success ratio, confirming its suitability for precise and consistent target-reaching tasks. Given these results, TD3 was selected for further evaluation in more complex tasks, as it showed better performance compared to DDPG in the initial tests.